# CytoNet: A Foundation Model for the Human Cerebral Cortex

### v1

@@ 45 Op5 Op6 Op7 Op8 Op9
Tabelle 3: List of areas from the Julich Brain Atlas (version 3.1, [116]) used in Figure 2. For brevity, pre- and postfixes to the area name are omitted (e.g., Area hOc1 (V1, 17, CalcS)“ is abbreviated as hOc1“). Indices of areas in the similarity matrices in Figure 2 are provided.
@@ 65 Id1 69 Id4 73 Id8 77 Ig3
Table 3 shows the list of areas from the Julich Brain Atlas (version 3.1, [116]) used for analyses presented in Figure 2. Table 2 provides the names for the 113 areas used in cytoarchitectonic classification, selected as the subset of atlas annotations available in at least four of the ten brains analyzed in this study.
@@ 7.2 Extended embedding analysis
Following common practice in self-supervised learning [118], CytoNet was trained with a shallow projection head appended to the backbone. We compared the latent space structure of features extracted from the backbone (our default), and after the shallow projection head of CytoNet-ViT (1M). Backbone features preserved brain identity (Figure 7A,C,E), whereas projected features showed minimal brain separation and instead aligned strongly with hemisphere (CHI 11577.88 → 28416.23) and atlas label (CHI 709.04 → 3648.73). Thus, the projection layers effectively suppressed brain-specific artifacts (e.g., staining, morphology, sectioning) and emphasized spatially consistent cytoarchitectonic organization. In the held-out brain, projections retained a global anterior–posterior structure but showed weaker differentiation by atlas label (CHI $77.63$ vs. $122.96\pm 12.17$) and hemisphere (CHI $27.85$ vs. $1633.94\pm 231.41$), reflecting reduced feature specificity.
@@
Complementing our 2D UMAP analyses (Figure 2), we visualized 3D UMAP embeddings colored by hemisphere, anterior-posterior and inferior-superior axes (Figure 8). The results show strong alignment between embedding locations and spatial locations, including distinct hemisphere-specific manifolds.
@@ 7.4 Extended scores for brain area classification and layer segmentation
7.5 Visualization of data-driven area classification through clustering
@@
Figure 9 visualizes the annotation-based clustering of areas Fp1 and Fp2 in B66 based on CytoNet-ViT (1M) features. Clusters were manually assigned to represent areas Fp1 or Fp2 based on visual inspection and compared to the reference annotations of the Julich Brain Atlas. The results show a strong alignment between annotations and cluster assignment, with an accuracy of $94.75\%$.

### v2

@@ 45 Op5 Op6 Op7 Op8 Op9
Tabelle 3: List of areas from the Julich Brain Atlas (version 3.1, [6]) used in Fig. 2. For brevity, pre- and postfixes to the area name are omitted (e.g., Area hOc1 (V1, 17, CalcS)“ is abbreviated as hOc1“). Indices of areas in the similarity matrices in Fig. 2 are provided.
@@ 65 Id1 69 Id4 73 Id8 77 Ig3
Table 3 shows the list of areas from the Julich Brain Atlas (version 3.1, [6]) used for analyses presented in Fig. 2. Table 2 provides the names for the 113 areas used in cytoarchitectonic classification, selected as the subset of atlas annotations available in at least four of the ten brains analyzed in this study.
@@ 11.2 Extended embedding analysis
Following common practice in self-supervised learning [19], CytoNet was trained with a shallow projection head appended to the backbone. We compared the latent space structure of features extracted from the backbone (our default), and after the shallow projection head of CytoNet-ViT (1M). Backbone features preserved brain identity (Fig. 7A,C,E), whereas projected features showed minimal brain separation and instead aligned strongly with hemisphere (CHI 11577.88 → 28416.23) and atlas label (CHI 709.04 → 3648.73). Thus, the projection layers effectively suppressed brain-specific artifacts (e.g., staining, morphology, sectioning) and emphasized spatially consistent cytoarchitectonic organization. In the held-out brain, projections retained a global anterior–posterior structure but showed weaker differentiation by atlas label (CHI $77.63$ vs. $122.96\pm 12.17$) and hemisphere (CHI $27.85$ vs. $1633.94\pm 231.41$), reflecting reduced feature specificity.
@@
Complementing our 2D UMAP analyses (Fig. 2), we visualized 3D UMAP embeddings colored by hemisphere, anterior-posterior and inferior-superior axes (Fig. 8). The results show strong alignment between embedding locations and spatial locations, including distinct hemisphere-specific manifolds.
@@ 11.4 Extended scores for brain area classification and layer segmentation
11.5 Visualization of data-driven mapping through clustering
@@
Fig. 9 visualizes the annotation-based clustering of areas Fp1 and Fp2 in B66 based on CytoNet-ViT (1M) features. Clusters were manually assigned to represent areas Fp1 or Fp2 based on visual inspection and compared to the reference annotations of the Julich Brain Atlas. The results show a strong alignment between annotations and cluster assignment, with an accuracy of $94.75\%$.